OpenStudy (anonymous):

The sum of three consecutive numbers is 99 what is the smallest of the three numbers?

OpenStudy (uglyturtle):

I would say 33.

OpenStudy (anonymous):

ok thank you so much :)

OpenStudy (nikato):

I don't think that'll work Let x= the smallest integer x+1= next integer x+2= the next,next integer Now you have the three consecutive integers and we know the sum is 99 So x+x+1+x+2=99 Can you solve for x?

OpenStudy (anonymous):

no haha I don't get it lol

OpenStudy (nikato):

Ok.. What part don't you get?

OpenStudy (anonymous):

where I have to slove for x

OpenStudy (nikato):

So you get the equation though? Ok you have x+x+1+x+2=99 Do you know how to combine like terms?

OpenStudy (anonymous):

eh not so much

OpenStudy (nikato):

Ok. I'll break it up for you. So lets just focus on the right side. I'm gonna rearrange so like terms are together x+x+x+1+2 So (x+x+x)+(1+2) What's x+x+x? And what's 1+2?

OpenStudy (anonymous):

ok so im thinking x means 1 and 1+2 =3 srry if im wrong

OpenStudy (nikato):

Yes. 1+2=3 And another way of writing x is 1x..so 1x+1x+1x=?

OpenStudy (anonymous):

3x

OpenStudy (nikato):

Yes. So now can you rewrite x+x+1+x+2=99 by simplifying the left side(combining like terms)

OpenStudy (anonymous):

ok I got stuck again srry

OpenStudy (nikato):

Rearranging You can x+x+x+1+2=99 And breaking it up like we did above, what will out left side look like now?

OpenStudy (anonymous):

um 9x?

OpenStudy (anonymous):

9x = 99 yea or no

OpenStudy (nikato):

3x+3=99 Do you see how I got this?

OpenStudy (anonymous):

oh yes I see now so that is the answer

OpenStudy (nikato):

Now you want solve for x. To start you can subtract 6 on both sides

OpenStudy (anonymous):

ok I know im suppose to subtract 6 from 99 but on the other side would I subtract 6 from 3x or 3 im thinking 3 but idk your the smart one here lol

OpenStudy (nikato):

3x+3=99 -3. -3 -------------- 3x=96. Remember 3-3=0 Does this make sense?

OpenStudy (anonymous):

yes keep going lol

OpenStudy (nikato):

Now can you divide both sides by 3? What do you get?

OpenStudy (nikato):

Oh my bad. It was suppose to say 3 not 6. Sorry if I was misleading you or something up there

OpenStudy (anonymous):

oh no it's ok

OpenStudy (campbell_st):

a simple method is divide the number by 3... them its the 2 integers either side of the answer. 99/3 = 33 so the numbers are 32, 33, 34

OpenStudy (kc_kennylau):

Let the smallest of them be x. The numbers will be x, (x+1), and (x+2) respectively.

OpenStudy (kc_kennylau):

Since their sum is 99, x+(x+1)+(x+2)=99

OpenStudy (kc_kennylau):

Understand this part?

OpenStudy (anonymous):

not really

OpenStudy (kc_kennylau):

If the smallest of them is x, then the next is x+1

OpenStudy (kc_kennylau):

Actually asking question is an essential part to learning

OpenStudy (kc_kennylau):

Therefore you should ask more questions

OpenStudy (kc_kennylau):

Where do you not get?

OpenStudy (kc_kennylau):

"The numbers will be x, (x+1), and (x+2) respectively." or "Since their sum is 99, x+(x+1)+(x+2)=99"?

OpenStudy (anonymous):

x=32

OpenStudy (kc_kennylau):

You said you couldn't get it, therefore I need to help you understand the concept

OpenStudy (anonymous):

nah im ok now thank you so much

OpenStudy (kc_kennylau):

no problem